OpenStudy (anonymous):

Solve this inequality. 7x + 2 > 30

OpenStudy (lgbasallote):

subtract 2 from both sides

OpenStudy (anonymous):

5x + 2 > 28?

OpenStudy (lgbasallote):

no..7x + 2 - 2 > 30 - 2

OpenStudy (anonymous):

7x + 0 > 28 ? or subtract 2 by 7 also?

OpenStudy (lgbasallote):

^that's right now divide both sides bu 7

OpenStudy (lgbasallote):

by*

OpenStudy (anonymous):

well you combine like terms when you subtract. you can only subtract from 2 and 30 since they are the only like terms

OpenStudy (anonymous):

i'm confused

OpenStudy (anonymous):

1x + 0 > 4

OpenStudy (anonymous):

???

OpenStudy (lgbasallote):

just write it simply as x > 4

OpenStudy (anonymous):

Yup 1x + 0 > 4 is long form 1x is just x cuzz any number times 1 is just that number aka 1*x=x any number + 0 is just that said number AKA x+0=x

OpenStudy (anonymous):

OOH and also the above properties are known as: 1 is known as multiplication ID and 0 as addition ID

OpenStudy (anonymous):

can you guys help me with this one -3x - 2 > 7 ? is it -3x - 2 > 5 = x - 2 > 3.2?